# Hot Jupiters are Inflated Primarily by Shallow Heating

### v1

@@ Abstract
The unexpectedly large radii of transiting hot Jupiters have led to many proposals for the physical mechanisms responsible for heating their interiors. While it has been shown that hot Jupiters reinflate as their host stars brighten due to heating deep in planetary interiors, young hot Jupiters also exhibit signs of delayed cooling possibly related to heating closer to their surfaces. To investigate this tension, we enhance our previously published hot Jupiter thermal evolution model by adding a parameter that allows for both deep heating and delayed cooling. We fit our thermal evolution models to a homogeneous, physically self-consistent catalog of accurate and precise hot Jupiter system properties in a hierarchical Bayesian framework. We find that hot Jupiters’ interior cooling rates are reduced on average by 95%–98% compared to simpler anomalous heating models. The most plausible explanation for this inference is substantial shallow heating just below their radiative–convective boundaries that enables reinflation with much less deep heating. Shallow heating by Ohmic dissipation and/or temperature advection are therefore important components of accurate models of hot Jupiter atmospheres, especially in circulation models. If hot Jupiters are inflated primarily by shallow heating as we propose, then we predict that their observed phase curve offsets should increase with temperature in the range $T_{\text{eq}}~\lesssim 1500~\text{K}$, peak in the range $1500~\text{K}~\lesssim~T_{\text{eq}}~\lesssim 1800~\text{K}$, and decrease in the range $T_{\text{eq}}~\gtrsim~1800~\text{K}$.
@@
journal: AAS Journalsfacilities: CDS, CTIO:2MASS, Exoplanet Archive, FLWO:2MASS, Gaia, GALEX, IRSA, NEOWISE, PS1, Skymapper, Sloan, WISEsoftware: astropy (AstropyI; AstropyII; AstropyIII), dustmaps (gre18), gaiadr3_zeropoint (lin21a), isochrones (mor15), matplotlib (hunter2007matplotlib), numpy (harris2020array), pandas (McKinney_2010; reback2020pandas), R (r25), scipy (jones2001scipy; 2020SciPy-NMeth), TOPCAT (tay05)
1 Introduction
@@
The radii of giant exoplanets that orbit very close to their host stars appear consistently larger than interior structure models would predict. These “hot Jupiters” have sufficiently small orbital separations that their strong instellation lengthens the atmospheric scale height and slows the loss of heat. It was expected that these planets would be somewhat inflated relative to the gas giants in our solar system. Indeed, even the first known transiting hot Jupiter, HD 209458 b, was found to be larger than Jupiter despite being nearly half its mass (Henry2000; Charbonneau2000). However, thermal evolution models of these planets struggled to match the substantially larger radii observed without introducing a powerful interior heat source (e.g., Bodenheimer2001; Guillot2002; Baraffe2003). As more hot Jupiters have been discovered via ground-based transit surveys like the Wide Angle Search for Planets (WASP; Pollacco2006; CollierCameron2007), Hungarian Automated Telescope Network (HATNet; Bakos2004; Bakos2007), Kilodegree Extremely Little Telescope (KELT; Pepper07), and Next Generation Transit Survey (NGTS; Wheatley18), as well as space-based transit surveys like the ESA’s CoRoT (Baglin06) mission and NASA’s Kepler (Borucki2010), K2 (Howell2014), and Transiting Exoplanet Survey Satellite (TESS; Ricker2014) missions, individual cases of larger-than-expected radii have developed into a clear trend for planets with $T_{\mathrm{eq}}>1000$ K (e.g., Miller2011; Demory2011; Thorngren2016). This discrepancy is now referred to as the hot Jupiter radius inflation problem [^1] See Thorngren2024 for a recent review of the topic., which asks why hot Jupiters remain large and internally hot for so long after they form.
@@
Statistical analyses of the hot Jupiter population have revealed some details about the nature of this problem. It was expected almost immediately after their initial discovery that hot Jupiter radius inflation should depend on some combination of a planet’s irradiation, mass, and age (e.g., Guillot2002; Burrows2004). As the population of hot Jupiters reached a size suitable for statistical analyses, stellar flux experienced by a hot Jupiter was identified as a far better predictor of a planet’s radius than its orbital period (Weiss2013). The effort to generate an empirical relation between planet radius and other independent variables was expanded upon by Thorngren2021, who found that incident flux and planet mass are strongly favored as predictors of planet radius, with a marginal additional dependence on stellar atmospheric metallicity.
@@
The observed correlation between incident flux and planet radius identified by Weiss2013 suggests that planets may grow in size as their parent stars brighten (Lopez2016). Grunblatt2016; Grunblatt2017 found evidence for such reinflation around post main-sequence stars. Additionally, Hartman2016 inferred similar reinflation with stellar evolution on the main sequence. Thorngren2021 reproduce this result and show it is a plausible outcome of planetary thermal evolution; however, this finding was complicated by a negative correlation between planet radius and age though. This was explained as a consequence of either more massive stars having shorter main sequence lifetimes than less massive stars or an additional mechanism that delays hot Jupiter cooling. The presence of more than one inflation mechanism complicates the picture, requiring more complex models to explain observed radius inflation.
@@
A primary differentiator between hot Jupiter heating mechanisms is their depth in the envelope, as this determines its effect on a planet’s radius evolution over time (Komacek2020). Some mechanisms, like Ohmic dissipation (Ohmic drag of particles driven by winds through a magnetic field; Batygin2010; Batygin2011; Perna2010a), compositional gradients (e.g., Chabrier2007), or fluid dynamical effects like turbulent mixing (Youdin2010) and longitudinal advection (Tremblin2017), are likely only able to deposit heat in the shallow interior. Such mechanisms make the planet’s temperature-pressure profile shallower, limiting the efficiency of outward heat escape and slowing radius contraction. However, as convection cannot move heat downwards, the only avenue for reinflation is the much slower conduction. Shallow heating therefore takes tens of Gyr to reinflate the planet — far too long to be observable.
@@
Other mechanisms like tidal dissipation (e.g., Bodenheimer2001; Leconte2010) or thermal tides (i.e., atmospheric temperature gradients causing tidal distortion in an attempt to reach equilibrium as suggested by Arras2009; Arras10; Socrates2013) may primarily occur deep in the interior. These mechanisms heat the planet throughout its interior, but do not slow the rate of heat escape from the interior. As such they can reinflate planets and keep them large, but would not produce the observed delayed cooling trends (e.g., Thorngren2021). To resolve this tension, we will model the observed hot Jupiter population as having both deep heating and delayed cooling.
@@
We will use a population-level analysis of the hot Jupiter population’s radius inflation to understand the relative contributions of deep heating and delayed cooling. By applying giant planet thermal evolution models to a hierarchical Bayesian inference framework, we will be able to use individual planet parameter inferences to infer a population-level value. This methodology has been used previously in similar circumstances: Thorngren2018 applied it to the population of hot Jupiters known at the time to extract the relation between heating efficiency and incident flux. Now we employ it with more advanced thermal evolution models.
@@
This analysis requires a homogeneous, physically self-consistent sample of planet and host star properties as an inhomogeneous sample of stellar masses, radii, and ages could bias the inferred cooling rates. In particular, precise ages are valuable for constraining the rate at which hot Jupiters cool. Age inference for mature exoplanet host stars remains difficult in the absence of asteroseismic or stellar mean density constraints from high-precision light curve data. However, our hierarchical Bayesian inference methodology mitigates the effects of limited information from a single system by aggregating results from many systems. This allows us to obtain a precise inference as attained by other population-level exoplanet age studies (e.g., Hamer2019; Schmidt24).
@@
In this article we will extract the population-level contributions of deep heating and delayed cooling to hot Jupiter radius inflation. We introduce our thermal evolution model in Section 2. We describe the construction of our homogeneous sample of hot Jupiter system properties in Section 3 and our application of these data to a hierarchical Bayesian inference framework in Section 4. We show our results and discuss their implications in Section 5, and summarize them Section 6.
2 Thermal Evolution Model
@@
Our planetary interiors are modeled similarly to Thorngren2023; we solve in one dimension the equations of hydrostatic equilibrium, mass conservation, and a suitable equation of state. We use Chabrier2021 for the H/He (with hydrogen and helium mass fractions as $X=0.754$ and $Y=0.246$ respectively), which in turn incorporates the Militzer2013 computations for the entropy. For the metals, we use a 50%/50% rock/ice mixture by mass from ANEOS (Thompson1990). These are placed in a fixed-mass $10~M_{\oplus}$ core; the remaining metal is mixed into the envelope via the additive volumes approximation.
In Thorngren2018, the thermal evolution was the sum of the intrinsic luminosity (calculated from the atmosphere models of Fortney2007) and an anomalous heating power $P_{a}(F)$ which was to be estimated:
@@
Here, $\sigma_{b}$ is the Stefan-Boltzmann constant and $T_{\mathrm{int}}$ is the intrinsic temperature. This model naturally allows the planet to reinflate, as it assumes all anomalous heating is able to efficiently heat the planet’s entire interior (e.g., Komacek2020). Conversely, cooling to equilibrium (defined as $dE/dt=0$) is very fast due to how strongly the intrinsic luminosity responds to changes in the adiabat entropy; this is shown as the dark blue “deep heating only” line in Fig. 1. By comparison, standard cooling with no anomalous heating is the light blue “cool giant” line, featuring relatively fast cooling and no reinflation.
@@
In this work we modify the model to allow for the delayed cooling and reinflation previously observed (e.g., Thorngren2021). Delayed cooling will be modeled with the intrinsic luminosity reduction factor $f_{\text{L}}$ that directly reduces the intrinsic luminosity of the planet. If included without any heating term, this modification would produce the light green “Delayed Cooling” line in Fig. 1 – note the large radii at young ages but a lack of reinflation. With both delayed cooling and anomalous heating, we have the dark green “Heated & Delayed” line, which features both large but decreasing radii at young ages as well as reinflation once the star ages and brightens sufficiently.
The effects of deep heating and delayed cooling have a compounding effect on the thermal and radius evolution of the planet. If $f_{\text{L}}=0.5$, then we would consequently only need half as much deep heating to maintain the $dE/dt=0$ equilibrium discussed in Thorngren2018 and Sarkis2020. To guarantee the same equilibrium state as previously observed, we scale both the intrinsic luminosity and the anomalous power by the same factor $f_{\text{L}}$. Thus our thermal evolution equation becomes
@@
$=2\pi R^{2}F\epsilon(F)f_{\text{L}}-4\pi R^{2}\sigma_{b}T_{\mathrm{int}}^{4}f_{\text{L}}+P_{\mathrm{rad}}.$ (3)
Following Thorngren2023, we have included radioactive heating $P_{\mathrm{rad}}$ for completeness; however, it is negligible for hot giants. We have also rewritten the anomalous heating as a fraction of the total instellation $\epsilon$, which we will refer to as the heating efficiency. Fig. 2 shows the radius evolution of an example planet for various values of $f_{\text{L}}$ and the median heating efficiency given by Thorngren2018. Only values of $f_{\text{L}}$ very close to zero can produce strong delayed cooling and begin to suppress reinflation.
Although we use an intrinsic luminosity reduction factor to parameterize this model, many models of hot Jupiter inflation mechanisms achieve this delay through shallow heating (see Sec. 1). Such effects could predominantly act by replacing outgoing heat from the interior rather than simply preventing its escape. In these cases, the intrinsic temperature observed in atmospheric mixing (e.g., Sing2024) would be the same as if the heating was all deep (no delayed cooling) despite the planet interior cooling more slowly. On the other hand, it is unclear how this would affect dynamo-dominated magnetic field strengths (see Yadav2017). Regardless, it is useful to note that our model is equivalent to shallow- and deep-heating powers of:
@@
Thus out-of-equilibrium shallow heating depends on $T_{\mathrm{int}}$ rather than $F$. This is necessary to achieve delayed cooling while still matching the observed radii – if we simply partitioned a flux-dependent heating into deep and shallow portions, we cannot reproduce both. While it seems more likely that shallow heating depends on both $T_{\mathrm{int}}$ and the intstellation $F$, such a model would be infeasible to fit from population radii. The model presented here should therefore be understood as a next-lowest-order approximation to the heating above that of Thorngren2018.
3 Host Star & Planet Property Inference
The inference of precise and accurate fundamental stellar parameters requires a precise parallax measurement. We therefore require Gaia Data Release (DR) 3 source_id identifiers for each hot Jupiter host in our inference sample [^2] For the details of Gaia DR3 and its data processing, see Gaia_mission2016; GaiaEDR3; GaiaDR3Frame; GaiaDR3Summary, GaiaEDR3Validation, lin21a; lin21b, GaiaEDR3XMatch; GaiaDR3XMatch, EDR3photometry, row21, tor21, and GaiaDR3Validation.. We select from the NASA Exoplanet Archive’s pscomppars table (ExoplanetArchive; Christiansen25; PSCompPars) as of July 22, 2025 all exoplanet systems available at the time with both a radius (pl_radj) and mass or minimum mass (pl_bmassj) measurement. We do not yet set limits on mass or radius as the stellar masses and radii we infer can cause the inferred planet masses and radii to differ from the reported composite parameters from the table. We assemble a list of Gaia DR3 source_id identifiers in three steps, following the procedure of Schmidt & Schlaufman (2025, submitted). We first use the Gaia DR2 source_id identifier provided by the NASA Exoplanet Archive to find the DR3 source_id using the dr2_neighbourhood table. If a NASA Exoplanet Archive entry lacks a Gaia DR2 source_id but has a TESS Input Catalog (TIC; TIC2019) ID, then we use the TIC ID to obtain a Gaia DR2 source_id and then follow the same procedure. If a NASA Exoplanet Archive entry lacks both a Gaia DR2 source_id and a TIC ID, then we use astroquery (astroquery) to query Simbad for the corresponding Gaia DR3 source_id corresponding to the system hostname. This procedure results in a sample of 1512 host stars amenable to our stellar property inference methodology.
We infer the fundamental and photospheric stellar parameters for the giant planet host stars in our sample using the isochrones (mor15) package to execute with MultiNest (fer08; fer09; fer19) a simultaneous Bayesian fit of the Modules for Experiments in Stellar Evolution (MESA; pax11; pax13; pax18; pax19; jer23) Isochrones & Stellar Tracks (MIST; dot16; cho16) isochrone grid to a curated collection of data for each star. Depending on the star, we fit the MIST grid to a subset of the following data:

### v2

# arXiv:2512.08932v2[astro-ph.EP] 02 Mar 2026

@@ Abstract
The unexpectedly large radii of transiting hot Jupiters have led to many proposals for the physical mechanisms responsible for heating their interiors. While it has been shown that hot Jupiters reinflate as their host stars brighten due to heating deep in planetary interiors, young hot Jupiters also exhibit signs of delayed cooling possibly related to heating closer to their surfaces. To investigate this ambiguity, we enhance our previously published hot Jupiter thermal evolution model by adding a parameter that allows for both deep heating and delayed cooling. We fit our thermal evolution models to a homogeneous, physically self-consistent catalog of accurate and precise hot Jupiter system properties in a hierarchical Bayesian framework. We find that hot Jupiters’ interior cooling rates are reduced on average by 95%–98% compared to simpler anomalous heating models. The most plausible explanation for this inference is substantial shallow heating just below their radiative–convective boundaries that enables reinflation with much weaker deep heating. Shallow heating by Ohmic dissipation and/or temperature advection are therefore important components of accurate models of hot Jupiter atmospheres, especially in circulation models. If hot Jupiters are inflated primarily by shallow heating as we propose, then we predict that atmospheric circulation-related observables should increase with temperature in the range $T_{\text{eq}}~\lesssim 1500~\text{K}$, peak in the range $1500~\text{K}~\lesssim~T_{\text{eq}}~\lesssim~1800~\text{K}$, and decrease in the range $T_{\text{eq}}~\gtrsim~1800~\text{K}$.
@@
journal: The Astrophysical Journalfacilities: CDS, CTIO:2MASS, Exoplanet Archive, FLWO:2MASS, Gaia, GALEX, IRSA, NEOWISE, PS1, Skymapper, Sloan, WISEsoftware: astropy (Astropy Collaboration et al., 2013, 2018, 2022), dustmaps (Green, 2018), gaiadr3_zeropoint (Lindegren et al., 2021a), isochrones (Morton, 2015), matplotlib (Hunter, 2007), numpy (Harris et al., 2020), pandas (McKinney, 2010; The pandas development team, 2020), R (R Core Team, 2025), scipy (Jones et al., 2024; Virtanen et al., 2020), TOPCAT (Taylor, 2005)
I Introduction
@@
The radii of giant exoplanets that orbit very close to their host stars appear consistently larger than interior structure models would predict. These “hot Jupiters” have sufficiently small orbital separations that their strong instellation lengthens the atmospheric scale height and slows the loss of heat. It was expected that these planets would be somewhat inflated relative to the gas giants in our solar system. Indeed, even the first known transiting hot Jupiter, HD 209458 b, was found to be larger than Jupiter despite being $\sim 70$% its mass (Henry et al., 2000; Charbonneau et al., 2000). However, thermal evolution models of these planets struggled to match the substantially larger radii observed without introducing a powerful interior heat source (e.g., Bodenheimer et al., 2001; Guillot and Showman, 2002; Baraffe et al., 2003). As more hot Jupiters have been discovered via ground-based transit surveys like the Wide Angle Search for Planets (WASP; Pollacco et al., 2006; Collier Cameron et al., 2007), Hungarian Automated Telescope Network (HATNet; Bakos et al., 2004, 2007), Kilodegree Extremely Little Telescope (KELT; Pepper et al., 2007), and Next Generation Transit Survey (NGTS; Wheatley et al., 2018), as well as space-based transit surveys like the ESA’s CoRoT (Baglin et al., 2006) mission and NASA’s Kepler (Borucki et al., 2010), K2 (Howell et al., 2014), and Transiting Exoplanet Survey Satellite (TESS; Ricker et al., 2014) missions, individual cases of larger-than-expected radii have developed into a clear trend for planets with $T_{\mathrm{eq}}>1000$ K (e.g., Miller and Fortney, 2011; Demory and Seager, 2011; Thorngren et al., 2016). This discrepancy is now referred to as the hot Jupiter radius inflation problem [^1] See Thorngren (2024) for a recent review of the topic., which asks why hot Jupiters remain large and internally hot for so long after they form.
@@
Statistical analyses of the hot Jupiter population have revealed some details about the nature of this problem. It was expected almost immediately after their initial discovery that hot Jupiter radius inflation should depend on some combination of a planet’s irradiation, mass, and age (e.g., Guillot and Showman, 2002; Burrows et al., 2004). As the population of hot Jupiters reached a size suitable for statistical analyses, hot Jupiter orbit-averaged equilibrium temperatures were found to correlate with the discrepancy between their measured radii and Bodenheimer et al. (2003) non-anomalously heated theoretical models (Laughlin et al., 2011). The stellar flux experienced by a hot Jupiter was subsequently identified as a far better predictor of a planet’s radius than its orbital period (Weiss et al., 2013). The effort to generate an empirical relation between planet radius and other independent variables was expanded upon by Thorngren et al. (2021), who found that incident flux and planet mass are strongly favored as predictors of planet radius, with a marginal additional dependence on stellar atmospheric metallicity.
@@
The observed correlation between incident flux and planet radius identified by Weiss et al. (2013) suggests that planets may grow in size as their parent stars brighten (Lopez and Fortney, 2016). Grunblatt et al. (2016, 2017) found evidence for such reinflation around post main-sequence stars. Additionally, Hartman et al. (2016) inferred similar reinflation with stellar evolution on the main sequence. Thorngren et al. (2021) reproduce this result and show it is a plausible outcome of planetary thermal evolution; however, this finding was complicated by a negative correlation between planet radius and age though. This was explained as a consequence of either more massive stars having shorter main sequence lifetimes than less massive stars or an additional mechanism that delays hot Jupiter cooling. The presence of more than one cause of unexpectedly large radii complicates the picture, requiring more complex models to explain observed radius inflation.
@@
A primary differentiator between hot Jupiter heating mechanisms is their depth in the envelope, as this determines its effect on a planet’s radius evolution over time (Komacek et al., 2020). Some mechanisms, like Ohmic dissipation (Lorentz forces on particles driven by winds through a magnetic field; Batygin and Stevenson, 2010; Batygin et al., 2011; Perna et al., 2010), compositional gradients (e.g., Chabrier and Baraffe, 2007), or fluid dynamical effects like turbulent mixing (Youdin and Mitchell, 2010) and radial potential temperature advection driven by superrotating longitudinal jets and other flows (Tremblin et al., 2017), are likely only able to deposit heat in the shallow interior. Such mechanisms make the planet’s temperature-pressure profile shallower, limiting the efficiency of outward heat escape and slowing radius contraction. However, as convection must cause a net upward heat flux, the only avenue for reinflation is the much slower conduction. Shallow heating therefore takes tens of Gyr to reinflate the planet — far too long to be observable.
@@
Other mechanisms like tidal dissipation (e.g., Bodenheimer et al., 2001; Leconte et al., 2010) or thermal tides (i.e., atmospheric temperature gradients causing tidal distortion in an attempt to reach equilibrium as suggested by Arras and Socrates, 2009, 2010; Socrates, 2013) may primarily occur deep in the interior. These mechanisms heat the planet throughout its interior, but do not slow the rate of heat escape from the interior. As such they can reinflate planets and keep them large, but would not produce the observed delayed cooling trends (e.g., Thorngren et al., 2021). To resolve this tension, we will model the observed hot Jupiter population as having both deep heating and delayed cooling [^2] To clarify, here we define hot Jupiter “heating models” as models that incorporate “mechanisms” that cause the planet’s radius to be “inflated,” i.e., larger than what would be expected without anomalous heating. While typically “inflation” would imply that the planet became larger over time, we use it as a catchall term for both “reinflation,” which follows the typical definition, and “delayed cooling,” which occurs in the opposite direction but still causes unexpectedly large radii..
@@
We will use a population-level analysis of the hot Jupiter population’s radius inflation to understand the relative contributions of deep heating and delayed cooling. By applying giant planet thermal evolution models to a hierarchical Bayesian inference framework, we will be able to use individual planet parameter inferences to infer a population-level value. This methodology has been used previously in similar circumstances: Thorngren and Fortney (2018) applied it to the population of hot Jupiters known at the time to extract the relation between heating efficiency and incident flux. Now we employ it with more advanced thermal evolution models.
@@
This analysis requires a homogeneous, physically self-consistent sample of planet and host star properties as an inhomogeneous sample of stellar masses, radii, and ages could bias the inferred cooling rates. In particular, precise ages are valuable for constraining the rate at which hot Jupiters cool. Age inference for mature exoplanet host stars remains difficult in the absence of asteroseismic or stellar mean density constraints from high-precision light curve data. However, our hierarchical Bayesian inference methodology mitigates the effects of limited information from a single system by aggregating results from many systems. This allows us to obtain a precise inference as attained by other population-level exoplanet age studies (e.g., Hamer and Schlaufman, 2019; Schmidt et al., 2024).
@@
In this article we will extract the population-level contributions of deep heating and delayed cooling to hot Jupiter radius inflation. We introduce our thermal evolution model in Section II. We describe the construction of our homogeneous sample of hot Jupiter system properties in Section III and our application of these data to a hierarchical Bayesian inference framework in Section IV. We show our results and discuss their implications in Section V, and summarize them Section VI.
II Thermal Evolution Model
@@
Our planetary interiors are modeled similarly to Thorngren et al. (2023); we solve in one dimension the equations of hydrostatic equilibrium, mass conservation, and a suitable equation of state. We use Chabrier and Debras (2021) for the H/He (with hydrogen and helium mass fractions as $X=0.754$ and $Y=0.246$ respectively), which in turn incorporates the Militzer and Hubbard (2013) computations for the entropy. For the metals, we use a 50%/50% rock/ice mixture by mass from ANEOS (Thompson, 1990). These are placed in a fixed-mass $10~M_{\oplus}$ core; the remaining metal is mixed into the envelope via the additive volumes approximation [^3] In cases where the planet is highly enriched in metals, the bulk of the metals are in the more compressible envelope rather than the core, leading to a smaller radius at a given metal enrichment and mass than a planet with a more massive core..
In Thorngren and Fortney (2018), the thermal evolution was the sum of the intrinsic luminosity (calculated from the atmosphere models of Fortney et al., 2007) and an anomalous heating power $P_{a}(F)$ which was to be estimated:
@@
Here, $\sigma_{b}$ is the Stefan-Boltzmann constant and $T_{\mathrm{int}}$ is the intrinsic temperature. This model naturally allows the planet to reinflate, as it assumes all anomalous heating is able to efficiently heat the planet’s entire interior (e.g., Komacek et al., 2020). Conversely, cooling to equilibrium (defined as $dE/dt=0$) is very fast due to how strongly the intrinsic luminosity responds to changes in the adiabat entropy; this is shown as the dark blue “deep heating only” line in Fig. 1. By comparison, standard cooling with no anomalous heating is the light blue “cool giant” line, featuring relatively fast cooling and no reinflation.
@@
In this work we modify the model to allow for the delayed cooling and reinflation previously inferred (e.g., Thorngren et al., 2021). Delayed cooling will be modeled with the intrinsic luminosity reduction factor $f_{\text{L}}$ that directly reduces the intrinsic luminosity of the planet. If included without any heating term, this modification would produce the light green “Delayed Cooling” line in Fig. 1 – note the large radii at young ages but a lack of reinflation. With both delayed cooling and anomalous heating, we have the dark green “Heated & Delayed” line, which features both large but decreasing radii at young ages as well as reinflation once the star ages and brightens sufficiently.
The effects of deep heating and delayed cooling have a compounding effect on the thermal and radius evolution of the planet. If $f_{\text{L}}=0.5$, then we would consequently only need half as much deep heating to maintain the $dE/dt=0$ equilibrium discussed in Thorngren and Fortney (2018) and Sarkis et al. (2020). To guarantee the same equilibrium state as previously observed, we scale both the intrinsic luminosity and the anomalous power by the same factor $f_{\text{L}}$. Thus our thermal evolution equation becomes
@@
$=\pi R^{2}F\epsilon(F)f_{\text{L}}-4\pi R^{2}\sigma_{b}T_{\mathrm{int}}^{4}f_{\text{L}}+P_{\mathrm{rad}}.$ (3)
Following Thorngren et al. (2023), we have included radioactive heating $P_{\mathrm{rad}}$ for completeness; however, it is negligible for hot giants. We have also rewritten the anomalous heating as a fraction of the total instellation $\epsilon$, which we will refer to as the heating efficiency. Fig. 2 shows the radius evolution of an example planet for various values of $f_{\text{L}}$ and the median heating efficiency given by Thorngren and Fortney (2018). Only values of $f_{\text{L}}$ very close to zero can produce strong delayed cooling and begin to suppress reinflation.
Although we use an intrinsic luminosity reduction factor to parameterize this model, many hot Jupiter heating models achieve this delay through shallow heating (see Sec. I). Such effects could predominantly act by replacing outgoing heat from the interior rather than simply preventing its escape. In these cases, the intrinsic temperature observed in atmospheric mixing (e.g., Sing et al., 2024) would be the same as if the heating was all deep (no delayed cooling) despite the planet interior cooling more slowly. On the other hand, it is unclear how this would affect dynamo-dominated magnetic field strengths (see Yadav and Thorngren, 2017). Regardless, it is useful to note that our model is equivalent to shallow- and deep-heating powers of:
@@
Thus out-of-equilibrium shallow heating depends on $T_{\mathrm{int}}$ rather than $F$. This is necessary to achieve delayed cooling while still matching the observed radii – if we simply partitioned a flux-dependent heating into deep and shallow portions, we cannot reproduce both. While it seems more likely that shallow heating depends on both $T_{\mathrm{int}}$ and the instellation $F$, such a model would be infeasible to fit from population radii. The model presented here should therefore be understood as a next-lowest-order approximation to the heating above that of Thorngren and Fortney (2018).
III Host Star & Planet Property Inference
The inference of precise and accurate fundamental stellar parameters requires a precise parallax measurement. We therefore require Gaia Data Release (DR) 3 source_id identifiers for each hot Jupiter host in our inference sample [^4] For the details of Gaia DR3 and its data processing, see Gaia Collaboration et al. (2016, 2021, 2022, 2023), Fabricius et al. (2021), Lindegren et al. (2021a, b), Marrese et al. (2021, 2022), Riello et al. (2021), Rowell et al. (2021), Torra et al. (2021), and Babusiaux et al. (2023).. We select from the NASA Exoplanet Archive’s pscomppars table (Akeson et al., 2013; Christiansen et al., 2025; NASA Exoplanet Archive, 2025a) as of July 22, 2025 all exoplanet systems available at the time with both a radius (pl_radj) and mass or minimum mass (pl_bmassj) measurement. We do not yet set limits on mass or radius as the stellar masses and radii we infer can cause the inferred planet masses and radii to differ from the reported composite parameters from the table. We assemble a list of Gaia DR3 source_id identifiers in three steps, following the procedure of Schmidt and Schlaufman (2026). We first use the Gaia DR2 source_id identifier provided by the NASA Exoplanet Archive to find the DR3 source_id using the dr2_neighbourhood table. If a NASA Exoplanet Archive entry lacks a Gaia DR2 source_id but has a TESS Input Catalog (TIC; Stassun et al., 2019) ID, then we use the TIC ID to obtain a Gaia DR2 source_id and then follow the same procedure. If a NASA Exoplanet Archive entry lacks both a Gaia DR2 source_id and a TIC ID, then we use astroquery (Ginsburg et al., 2019) to query Simbad for the corresponding Gaia DR3 source_id corresponding to the system hostname. This procedure results in a sample of 1512 host stars amenable to our stellar property inference methodology.
We infer the fundamental and photospheric stellar parameters for the giant planet host stars in our sample using the isochrones (Morton, 2015) package to execute with MultiNest (Feroz and Hobson, 2008; Feroz et al., 2009, 2019) a simultaneous Bayesian fit of the Modules for Experiments in Stellar Evolution (MESA; Paxton et al., 2011, 2013, 2018, 2019; Jermyn et al., 2023) Isochrones & Stellar Tracks (MIST; Dotter, 2016; Choi et al., 2016) isochrone grid to a curated collection of data for each star. Depending on the star, we fit the MIST grid to a subset of the following data: